If 1/f=1/u +1/v

Were u =10.5+-0.1cm and V =85+-0.1cm find minimum error in f


1
Expert's answer
2021-04-13T09:10:45-0400

To be given in question

"u=" 10.5 cm

"\Delta u" =-0.1 cm

"v=" 85 cm

"\Delta v=" -0.1 cm 

To be asked in question

"{\Delta f}=f(\frac{\Delta u}{u}+\frac{\Delta v}{v})"

We know that

"\frac{1}{f}=\frac{1}{u}+\frac{1}{u}"

Put value then

"f=9.3455cm"

"\frac{\Delta f}{f}=\frac{\Delta u}{u}+\frac{\Delta v}{v}"

"{\Delta f}=f(\frac{\Delta u}{u}+\frac{\Delta v}{v})"

Put value

"\Delta f=-9.3455(\frac {-0.1}{10.5}+\frac {-0.1}{85})"

"\Delta f=-0.0999cm"